Grade 11Mechanics

a body of mass m thrown horizontally with a velocity v from the top of a tower of height h touches the ground at a distance 250 m from the foot of tower . Another body of mass 2m thrown horizontally with half the velocity from a tower of height h will touch the ground at a distance

liopk
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Dear Student,
From Distance = speed x time we know that,
R=v.t
For both the cases, h= -0+gt2/2
=>t=√(2h/g)
 
R for the second ball = v/2.t = R/2 = 125m
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)